OpenStudy (anonymous):

If the midpoint between (8, y) and (-11, 6) is (-1.5, 5), find the value of y how do i get this formula 5 = ( 6 + y)/2

myininaya (myininaya):

the midpoint=(average of the x numbers of the endpoints, average of the y numbers of the end endpoints)

myininaya (myininaya):

you have 5=(y+6)/2 and you have -1.5=(8+(-11))/2

myininaya (myininaya):

the second equation is true you only need to solve 5=(y+6)/2

OpenStudy (anonymous):

how whould ik what to put where

myininaya (myininaya):

what do you mean?

myininaya (myininaya):

the y numbers of the endpoints are y and 6 the average of y and 6 is given by (y+6)/2 where this number is also suppose to be 5

OpenStudy (anonymous):

if i have this If the midpoint between (18, y) and (20, -15) is (19, -5), find the value of y. it whould be in this form 18=20(+19)/2

myininaya (myininaya):

 and you do the same thing with the y's..